​(Related to Checkpoint 14.2 and Checkpoint​ 14.3) ​ (Cost of common​ equity)  The common stock for the Hetterbrand Corporation sells for ​$59.82​, and the last dividend paid was ​$2.23. Five years ago the firm paid ​$1.93 per​ share, and dividends are expected to grow at the same annual rate in the future as they did over the past five years. a. What is the estimated cost of common equity to the firm using the dividend growth​ model? b. ​ Hetterbrand's CFO has asked his financial analyst to estimate the​ firm's cost of common equity using the CAPM as a way of validating the earlier calculations. The​ risk-free rate of interest is currently 4.8 ​percent, the market risk premium is estimated to be 5.4 ​percent, and​ Hetterbrand's beta is 0.75. What is your estimate of the​ firm's cost of common equity using this​ method? a.  The estimated cost of common equity to the firm using the dividend growth model is nothing​%. ​(Round to two decimal​ places.)
Answer #1

A: Cost of equity = Dividend expected/ Price + growth

Find growth rate using the FV formula

FV = PV*(1+g)^n

2.23=1.93*(1+g)^5

G = 2.93%

Cost of equity = 2.23*102.93%/59.82 + 2.93%

=6.77%

B: Cost of equity using CAPM = Rf+ Beta*Rpm

= 4.8%+ 0.75*5.4%

=8.85%